calculate the mass of urea required inmaking 2.5kg of 0.5molal aqueous solution?
 

Grade:11

1 Answers

Mredul Jain
7 years ago
Let mass of urea added =xkg=> mass of water added =(2.5-x)kg=> (1000*x/(60))*1/(2.5-x)= 0.5=> x= 75/31 gx=2.42 g